12g of an alkaline earth metal gave 14.8g of its nitride.Atomic wt of that metal is

Grade:12

4 Answers

Ramesh V
14 years ago

let the metal be M with atmic wt. be x

so 3M + N2 ---->M3N2

      3x                   3x+28

 so,  12/3x   = 14.8/(3x+28)

which in solving for x gives 40

the metal is calcium and its nitride is Ca3N2

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it
here and we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE
preparation. All the best.

Regards,
Ramesh
IIT Kgp - 05 batch

mohd atif chaudhary
5 years ago
Mass of metal is 12g and Mass of metal nitride is 14.8gm so Mass of nitride ion is 2.8 gm.
By using equivalence law: Mass of metal / Mass of Nitride ion =Equivalent Mass of metal/14 
12/2.8=3x/14
x=20 as x is equivalent weight by multiplying it with 2 which is valency of metal ion. We get 40. 
Before x is 3 is used which is valency of Nitride and 14 is Mass of N .
 
mohd atif chaudhary
5 years ago
Mass of metal is 12g and Mass of metal nitride is 14.8gm so Mass of nitride ion is 2.8 gm.By using equivalence law: Mass of metal / Mass of Nitride ion =Equivalent Mass of metal/14 12/2.8=3x/14x=20 as x is equivalent weight by multiplying it with 2 which is valency of metal ion. We get 40. Before x is 3 is used which is valency of Nitride and 14 is Mass of N .
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Let the metal be M with atmic wt. be x
so 3M + N2 ---->M3N2
     3x                3x+28
 
so,  12/3x   = 14.8/(3x+28)
Hence, x = 40
Hence, the metal is calcium and its nitride is Ca3N2
And the molar mass of metal (Ca) is 40
 
Thanks and regards,
Kushagra